# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

### v1

@@ Abstract
We make three contributions to conformal prediction. First, we propose fuzzy conformal confidence sets that offer a degree of exclusion, generalizing beyond the binary inclusion/exclusion offered by classical confidence sets. We connect fuzzy confidence sets to e-values to show this degree of exclusion is equivalent to an exclusion at different confidence levels, capturing precisely what e-values bring to conformal prediction. We show that a fuzzy confidence set is a predictive distribution with a more appropriate error guarantee. Second, we derive optimal conformal confidence sets by interpreting the minimization of the expected measure of the confidence set as an optimal testing problem against a particular alternative. We use this to characterize exactly in what sense traditional conformal prediction is optimal. Third, we generalize the inheritance of guarantees by subsequent minimax decisions from confidence sets to fuzzy confidence sets. All our results generalize beyond the exchangeable conformal setting to prediction sets for arbitrary models. In particular, we find that any valid test (e-value) for a hypothesis automatically defines a (fuzzy) prediction confidence set.
Keywords: Conformal prediction, prediction sets, e-values, fuzzy confidence sets, certified decisions.
@@ 1.1 Traditional conformal prediction
Suppose we have $n$ exchangeable observations $Z^{n}=(Z_{1},\dots,Z_{n})$. Abstracting away from regressors, the goal of conformal prediction is to use these $n$ observations to construct a confidence set $C_{\alpha}^{Z^{n}}$ that will cover the next observation with a probability of at least $1-\alpha$,
@@
To construct such a confidence set, the idea behind conformal prediction is to plug-in hypothetical values $z$ for $Z_{n+1}$ and test whether $(Z_{1},\dots,Z_{n},z)$ is exchangeable at level $\alpha$. Repeating this exercise for each plug-in value $z$ in the sample space, the confidence set is formed by collecting the values $z$ for which the hypothesis is not rejected at level $\alpha$:
@@
Figure 1: A traditional and fuzzy conformal confidence set. The solid line represents a traditional level $\alpha=0.05$ confidence set: the points for which this line is zero form the confidence set, and the points at which it equals 11 its complement. The dashed line represents a fuzzy $[0,1]$-valued confidence set, offering a degree of exclusion at each point.
@@ 1.2 Fuzzy conformal prediction
A downside of a traditional conformal confidence set is that it is binary in nature: a sample point $z$ either falls inside the confidence set or not. To solve this, we propose fuzzy conformal confidence sets.
For each point $z\in\mathcal{Z}$, a traditional confidence set may be reinterpreted as a function $C_{\alpha}^{Z^{n}}:\mathcal{Z}\to\{0,1\}$, which outputs an inclusion (0) or exclusion (1) for each point $z\in\mathcal{Z}$. A fuzzy confidence set $\widetilde{C}_{\alpha}^{Z^{n}}:\mathcal{Z}\to[0,1]$ generalizes beyond the binary inclusion-exclusion to a degree of exclusion, in the interval $[0,1]$. We illustrate this in Figure 1, where the solid line represents a traditional confidence set and the dashed line a fuzzy confidence set.
For a fuzzy confidence set, the traditional coverage guarantee (1) generalizes to the guarantee that the degree of exclusion of $Z_{n+1}$ is bounded in expectation by $\alpha$:
@@
1.3 E-values and fuzzy confidence
An apparent downside of a fuzzy confidence set is that the ‘degree of exclusion of $z$ at level $\alpha$’ is difficult to interpret. We overcome this by merging the degree of exclusion into the significance level, allowing different points $z$ to be excluded at different significance levels. This relies on the recently established equivalence between fuzzy tests and ‘e-values’ (Koning, 2025b).
To facilitate this, we rescale the confidence set by dividing it by $\alpha$:
@@ 1.3 E-values and fuzzy confidence
Due to the rescaling, $\mathcal{E}^{Z^{n}}(z)=1/\alpha$ now corresponds to a full exclusion at level $\alpha$. More generally, we show that the value of a fuzzy confidence set $\mathcal{E}^{Z^{n}}(z)$ at each point $z$ may be directly interpreted as (the reciprocal of) the smallest significance level at which $z$ is excluded, under a generalized Type I error guarantee (Grünwald, 2023a; Koning, 2024). This replaces the ‘degree of exclusion of a point $z$ at level $\alpha$’ by an ‘exclusion at level $\alpha^{\prime}=1/\mathcal{E}^{Z^{n}}(z)$’. Because of the rescaling, the guarantee (2) must be reformulated to
@@
which also enables $0,\infty]-$valued fuzzy confidence sets.
We illustrate fuzzy confidence sets on this evidence scale in Figure 2, which is a rescaled version of Figure 1, now also including a $0,\infty]-$valued fuzzy confidence set. Here, we see that the traditional confidence set only offers an exclusion at level $1/20=0.05$, whereas the fuzzy confidence sets offer exclusions at various significance levels for different points in the sample space. Compared to the traditional confidence set, the unbounded fuzzy confidence set expresses very strong evidence against large values of $z_{n+1}$. The bounded version is capped at evidence corresponding to a rejection at level $1/20=0.05$, but trades this against more evidence for small values of $z_{n+1}$ compared to the unbounded version.
Figure 2: A traditional and fuzzy conformal confidence sets on the evidence scale. The solid line represents a traditional level $\alpha=0.05$ confidence set, the dashed line represents a fuzzy $0,1/\alpha]-$valued confidence set, and the dotted line a fuzzy $0,\infty]-$valued confidence set. An evidence value of $\mathcal{E}^{Z^{n}}(z)=e$ corresponds to an exclusion at significance level $1/e$.
1.4 Optimal prediction sets
Besides generalizing to fuzzy confidence sets, we also contribute to classical non-fuzzy conformal prediction by studying its optimality. In fact, our optimality results generalize beyond exchangeability to prediction confidence sets for arbitrary models.
The key to unlocking our optimality results is to show that a prediction confidence set is equivalent to a single hypothesis test, which is valid for a model $\mathcal{P}$ if and only if the test is valid for the hypothesis $\mathcal{P}$. This means that any test that is valid for a hypothesis $\mathcal{P}$ automatically defines a prediction confidence set.
Next, we use the equivalence between a prediction confidence set and a test to study optimal confidence sets. In particular, we show that a confidence set is of minimal expected size under some measure $\mu$,
@@ 1.4 Optimal prediction sets
if and only if the corresponding test is most powerful against some alternative hypothesis that incorporates $\mu$. More generally, this measure may be a kernel $\mu^{|Z^{n}}$ that is allowed to depend on $Z^{n}$.
In Section 4, we specialize the optimality to the classical conformal prediction setting where $\mathcal{P}$ is the exchangeable model to derive optimal confidence sets for conformal prediction. We use this to show in what sense classical conformal prediction is optimal, given a choice of conformity score.
We show these results generalize to fuzzy confidence sets, which we show are equivalent to e-values. Optimal fuzzy confidence sets may then be derived from e-values that are optimal against a particular utility function $U$. Classical optimal confidence sets are recovered when using the ‘Neyman–Pearson utility function’ $U:x\mapsto x\wedge 1/\alpha$, which correspond to a Neyman–Pearson most powerful test. This Neyman–Pearson utility function expresses a linear preference up to $1/\alpha$, and no interest in evidence beyond a rejection at the pre-specified level $\alpha$. Other utility functions may be used to express other preferences, which subsequently shape the fuzzy confidence set.
@@ 1.5 Certifying subsequent decisions
As a final contribution, treated in Section 5, we explain how a fuzzy confidence set can be used in subsequent decision making. Here, we show how the coverage guarantee offered by a fuzzy confidence set may be inherited by subsequent minimax decisions. This generalizes ideas of Kiyani et al. (2025) from traditional conformal confidence sets to fuzzy confidence sets and general prediction confidence sets.
@@
Most of our findings here build on results of Grünwald (2023b) and Andrews and Chen (2025) to conformal prediction. This relies on the surprising (to us) observation that a fuzzy confidence set is equivalent to (the reciprocal of) an E-posterior; an object recently studied by Grünwald (2023b). This shows that the ‘E-posterior’ is deeply connected to classical hypothesis testing. Furthermore, our perspective unifies the notions of P-certification and E-certification that were recently developed by Andrews and Chen (2025).
1.6 Related literature
@@
To the best of our knowledge, fuzzy confidence sets were first explicitly proposed by Geyer and Meeden (2005) for confidence sets on parameter spaces. Compared to their work, we deliver three innovations. First, Geyer and Meeden (2005) stick to a classical Neyman–Pearson-style power objective, limiting the use of fuzzy confidence sets to discrete data. We generalize to different expected-utility-based power objectives, leading to properly fuzzy confidence sets even in non-discrete data. Second, we show that a ‘degree of inclusion at level $\alpha$’ is equivalent to an exclusion at a different significance level under a generalized Type I error guarantee. Third, we consider fuzzy prediction confidence sets, which are on the sample space instead of on a parameter space.
Vovk (2024) states that precursors of conformal prediction actually relied on e-values (under a different name). He attributes their demise in conformal prediction to the fact that they found obtaining the confidence guarantee more natural with p-values. This is in harmony with our finding that e-values are equivalent to fuzzy confidence sets, showing that e-values play no role in classical (non-fuzzy) confidence sets. Moreover, Vovk (2024) cites the advantage of p-values for constructing conformal predictive distributions. In Remark 2, we show that conformal predictive distributions are (the reciprocal of) defective fuzzy confidence sets. Equivalently, we find that fuzzy conformal confidence sets are (the reciprocal of) conformal predictive distributions that offer an additional desirable joint guarantee, which may be inherited by subsequent decisions.
To the best of our knowledge, the first explicit connection between conformal prediction and e-values is made in the first ArXiv version of Wang and Ramdas (2022) (but omitted in the journal version), where they briefly mention conformal prediction as a potential application of a ‘softrank’ e-value they derive for exchangeability. Koning (2025a) independently derived a more general class of e-values of the form
$\frac{T(Z^{n+1})}{1/(n+1)!\sum_{\pi}T(\pi Z^{n+1})},$ (3)
where $\pi Z^{n+1}$ denotes a permutation $\pi$ of the elements of the tuple $Z^{n+1}$, which nests softrank e-value for a particular choice of the non-negative statistic $T$. Moreover, he treats conformal prediction with e-values in two brief examples, highlighting the potential for post-hoc selection of $\alpha$ (Grünwald, 2023a; Koning, 2024). Soon after, Balinsky and Balinsky (2024) independently discovered (3), using such e-values as a tool to construct traditional conformal confidence sets. More recently, Gauthier et al. (2025) reconnected e-values to the post-hoc selection of $\alpha$, also focusing on e-values of the form (3).
We make several contributions to this line of literature. First, we use the equivalence between e-values and fuzzy testing derived by Koning (2025b) to show that fuzzy confidence sets are precisely what e-values offer to conformal prediction. Second, while such a fuzzy confidence set permits the subsequent post-hoc selection of a single confidence set, we explicitly recommend not to do so, and propose to instead report the entire fuzzy confidence set. Third, and perhaps most important, preceding applications of e-values in conformal prediction rely on e-values of the form (3), but provide little to no guidance on how to select the statistic $T$. This is problematic, as any admissible e-value is of this form (Koning, 2025a). We instead derive optimal e-values for conformal prediction against alternatives that yield ‘small’ fuzzy conformal confidence sets. Finally, the preceding work considers e-values and conformal prediction within the classical exchangeability model, while our work generalizes to prediction confidence sets for arbitrary models.
2 Prediction confidence sets
2.1 Coverage beyond exchangeability
@@
Let $\mathcal{Z}$ denote the sample space of a single observation, and $\mathcal{Z}^{n+1}$ the joint sample space of $(n+1)$ observations. A prediction confidence set is classically defined as a set $C_{\alpha}^{Z^{n}}$ with the level $\alpha>0$ coverage guarantee
@@
where $\mathcal{P}$ is some model (collection of distributions) on $\mathcal{Z}^{n+1}$. We say that such a confidence set is valid for $\mathcal{P}$. Conformal prediction is the special case where the model $\mathcal{P}$ is the collection $\mathcal{P}^{\textnormal{exch.}}$ of exchangeable distributions on $\mathcal{Z}^{n+1}$, but we delay specializing to exchangeability in Section 4.
Note that the probability in (4) is over both $Z^{n}$ and $Z_{n+1}$, and so over the entire tuple $Z^{n+1}$. This means that the confidence set may be interpreted as a function of $Z^{n}$. For notational brevity, we write $C_{\alpha}$ to denote this function, and $C_{\alpha}^{z^{n}}$ for the function evaluated at $Z^{n}=z^{n}\in\mathcal{Z}^{n}$.
2.2 The non-coverage event as a test
Our first insight is that a prediction confidence set $C_{\alpha}$ is equivalent to a single test $\phi:\mathcal{Z}^{n+1}\to\{0,1\}$. Moreover, the confidence set is valid for a model $\mathcal{P}$ if and only if the corresponding test is valid for the hypothesis $\mathcal{P}$. A consequence of this result is that any test $\phi$ that is valid for some hypothesis $\mathcal{P}$ automatically produces a prediction confidence set for $\mathcal{P}$.
Due to the importance of this insight, we capture it in Theorem 1. We cover its proof in the main text, as it is simple and constructive.

### v2

# Fuzzy Prediction Sets: Conformal Prediction with E-values

@@ Abstract
We make three contributions to conformal prediction. First, we propose fuzzy conformal prediction sets that offer a degree of exclusion, generalizing beyond the binary inclusion/exclusion offered by classical prediction sets. We connect fuzzy prediction sets to e-values to show this degree of exclusion is equivalent to an exclusion at different confidence levels, capturing precisely what e-values bring to conformal prediction. We show that a fuzzy prediction set is a predictive distribution with an arguably more appropriate error guarantee. Second, we derive optimal conformal prediction sets by interpreting the minimization of the expected measure of a prediction set as an optimal testing problem against a particular alternative. We use this to characterize exactly in what sense traditional conformal prediction is optimal, and show how this may generally be used to construct optimal (fuzzy) prediction sets. Third, we generalize the inheritance of guarantees by subsequent minimax decisions from prediction sets to fuzzy prediction sets. All results generalize beyond the conformal setting to prediction sets for arbitrary models. In particular, we find that constructing a (fuzzy) prediction set for a model is equivalent to constructing a test (e-value) for that model as a hypothesis.
Keywords: Conformal prediction, prediction set, e-value, fuzzy confidence set, certified decision.
@@ 1.1 Traditional conformal prediction
Suppose we have $n$ exchangeable observations $Z^{n}=(Z_{1},\dots,Z_{n})$. Abstracting away from covariates, the goal of conformal prediction is to use these $n$ observations to construct a prediction set $C_{\alpha}^{Z^{n}}$ that will cover the next observation with a probability of at least $1-\alpha$,
@@
To construct such a prediction set, the idea behind conformal prediction is to plug-in hypothetical values $z$ for $Z_{n+1}$ and test whether $(Z_{1},\dots,Z_{n},z)$ is exchangeable at level $\alpha$. Repeating this exercise for each plug-in value $z$ in the sample space, the prediction set is formed by collecting the values $z$ for which the hypothesis is not rejected at $\alpha$:
@@
Figure 1: A traditional and fuzzy conformal prediction set. The solid line represents a traditional level $\alpha=0.05$ prediction set: the points $z_{n+1}$ at which this line is zero form the prediction set, and the points at which it equals 11 its complement. The dashed line represents a fuzzy $[0,1]$-valued prediction set, offering a degree of exclusion at each point.
@@ 1.2 Fuzzy conformal prediction
A downside of a traditional conformal prediction set is that it is binary in nature: a sample point $z$ either falls inside the set or not. To tackle this, we propose fuzzy conformal prediction sets.
For each point $z\in\mathcal{Z}$, a traditional prediction set may be reinterpreted as a function $C_{\alpha}^{Z^{n}}:\mathcal{Z}\to\{0,1\}$, which outputs an inclusion (0) or exclusion (1) for each point $z\in\mathcal{Z}$. A fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}:\mathcal{Z}\to[0,1]$ generalizes beyond the binary inclusion-exclusion to a degree of exclusion, in the interval $[0,1]$. We illustrate this in Figure 1, where the solid line represents a traditional prediction set and the dashed line a fuzzy prediction set.
For a fuzzy prediction set, the traditional coverage guarantee (1) generalizes to the guarantee that the degree of exclusion of $Z_{n+1}$ is bounded in expectation by $\alpha$:
@@
1.3 E-values and fuzzy prediction
An apparent downside of a fuzzy prediction set is that the ‘degree of exclusion of $z$ at level $\alpha$’ is difficult to interpret. We overcome this by merging the degree of exclusion into the significance level, allowing different points $z$ to be excluded at different significance levels. This relies on the recently established equivalence between fuzzy tests and ‘e-values’ (koning2025continuoustestingunifyingtests).
To facilitate this, we rescale the prediction set by dividing it by $\alpha$:
@@ 1.3 E-values and fuzzy prediction
Due to the rescaling, $\mathcal{E}^{Z^{n}}(z)=1/\alpha$ now corresponds to a full exclusion at level $\alpha$. More generally, we show that the value of a fuzzy prediction set $\mathcal{E}^{Z^{n}}(z)$ at any point $z$ may be directly interpreted as (the reciprocal of) the smallest significance level at which $z$ is excluded, under a generalized Type I error guarantee (grunwald2022beyond; koning2024post). This replaces the ‘degree of exclusion of a point $z$ at level $\alpha$’ by an ‘exclusion at level $\alpha^{\prime}=1/\mathcal{E}^{Z^{n}}(z)$’. Because of the rescaling, the guarantee (2) must be reformulated to
@@
which also enables $0,\infty]-$valued fuzzy prediction sets.
We illustrate fuzzy prediction sets on this evidence scale in Figure 2, which is a rescaled version of Figure 1 that also includes a $0,\infty]-$valued fuzzy prediction set. Here, we see that the traditional prediction set only offers an exclusion at level $1/20=0.05$, whereas the fuzzy prediction sets offer exclusions at various significance levels for different points in the sample space. Compared to the traditional prediction set, the unbounded fuzzy prediction set expresses very strong evidence against large values of $z_{n+1}$. The bounded version is capped at evidence corresponding to a rejection at level $1/20=0.05$, but trades this against more evidence for small values of $z_{n+1}$ compared to the unbounded version.
We show that fuzzy prediction sets may easily be merged, by averaging their pointwise evidence under arbitrary dependence, or multiplying under independence. Applying this averaging to non-fuzzy prediction sets generally produces fuzzy prediction sets.
Figure 2: One traditional and two fuzzy prediction sets on the evidence scale. The solid line represents a traditional level $\alpha=0.05$ prediction set, the dashed line represents a fuzzy $0,1/\alpha]-$valued prediction set, and the dotted line a fuzzy $0,\infty]-$valued prediction set. An evidence value of $\mathcal{E}^{Z^{n}}(z)=a$ corresponds to an exclusion at significance level $1/a$.
1.4 Optimal (fuzzy) prediction sets
Besides generalizing to fuzzy prediction sets, we also contribute to classical non-fuzzy conformal prediction by studying its optimality. In fact, our optimality results generalize beyond exchangeability to prediction sets for arbitrary models.
The key to unlocking our optimality results is to show that a prediction set for a model $\mathcal{P}$’ is equivalent to a single test for the hypothesis $\mathcal{P}$. We subsequently link optimal prediction sets to optimal testing theory. In particular, we show that a prediction set is of minimal expected size under some measure $\mu$,
@@ 1.4 Optimal (fuzzy) prediction sets
if and only if the corresponding test is most powerful against some alternative hypothesis that incorporates $\mu$.
In Section 5, we specialize the optimality to the classical conformal prediction setting where $\mathcal{P}$ is the exchangeable model to derive optimal prediction sets for conformal prediction. We use this to show in what sense classical conformal prediction is optimal, given a choice of conformity score.
These results generalize to fuzzy prediction sets, which are equivalent to e-values. We connect optimal fuzzy prediction sets to expected-utility optimal e-values koning2025continuoustestingunifyingtests. Classical conformal prediction is recovered as a special case, for the ‘Neyman–Pearson utility function’ $U^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, when the e-value corresponds to the Neyman–Pearson most powerful test at level $\alpha$. Other utility functions may be used to express different preferences regarding the value of different amounts of evidence, which subsequently shapes the fuzzy prediction set.
@@ 1.5 Certifying subsequent decisions
In Section 6, we explain how a fuzzy prediction set can be used in subsequent decision making. Here, we show how the coverage guarantee offered by a fuzzy prediction set may be inherited by subsequent minimax decisions. This generalizes ideas of kiyani2025decision from traditional conformal prediction sets to fuzzy conformal prediction sets and general prediction sets.
@@
Most of our findings here build on and extend results of grunwald2023posterior and andrews2025certified to conformal prediction. This relies on the surprising (to us) observation that a fuzzy confidence set is equivalent to (the reciprocal of) an E-posterior; an object recently studied by grunwald2023posterior. This shows that the ‘E-posterior’ is deeply connected to classical hypothesis testing. Furthermore, our perspective unifies the notions of P-certification and E-certification that were recently developed by andrews2025certified.
1.6 Covariates and application
In Section 7, we show how our results extend to the setting of covariates. In the covariate setting, the tuple $Z^{n+1}$ can be decomposed into $Z^{n+1}=(X^{n+1},Y^{n+1})$, and we measure the covariate $X_{n+1}$ alongside $Z^{n}$. The goal is then to obtain a prediction set for $Y_{n+1}$. Although the use of covariates is standard in conformal prediction, we defer their inclusion to this later section as they add considerable notational overhead and are not essential to our main results.
We apply our methodology to a classical character-recognition problem in Section 8. Here, we showcase what fuzzy prediction sets offer over classical prediction sets, and study how different utility functions shape the fuzzy prediction sets.
2 Related literature
@@
To the best of our knowledge, fuzzy confidence sets were first explicitly proposed by geyer2005fuzzy for confidence sets on parameter spaces. Compared to their work, we deliver several innovations. First, while conceptually the same type of object, we stress that our work is on fuzzy prediction sets, which may be viewed as confidence sets on a sample space instead of on a parameter space. Second, geyer2005fuzzy present fuzzy confidence sets as a solution to under-coverage of classical Neyman–Pearson optimal confidence sets for discrete data in small samples. We advocate for a much wider use, beyond discrete data, by using a generalization of the Neyman–Pearson framework from the recent e-value literature. Third, we show that a fuzzy ‘degree of exclusion at level $\alpha$’ is equivalent to an exclusion at a different significance level under a generalized Type I error guarantee. We believe this may make fuzzy prediction/confidence sets more palatable, as this makes them easier to interpret and work with. Finally, we also attempt to show the relevance of fuzzy prediction sets by showing how they may be used in subsequent decision making, further motivating their relevance over non-fuzzy prediction sets.
As the literature on conformal prediction and e-values remains fairly small, we attempt to give a complete overview. According to vovk2024conformal, precursors of conformal prediction (Gammerman1998; Vovk1999) actually relied on e-values (under a different name). He attributes their demise in conformal prediction to the fact that prediction sets were more naturally obtained by using p-values. This is in line with our finding that e-values are equivalent to fuzzy prediction sets, which implies that e-values play no fundamental role in classical (non-fuzzy) prediction sets.
The first explicit connection between modern conformal prediction and e-values appears in the 2020 arXiv version of vovk2024conformal, where he exploits the merging properties of e-values to enable cross-conformal prediction by averaging over data splits. Since then, several other works have used e-values as a pragmatic tool to construct conformal prediction sets. For example, in the context of conformal selection, lee2024boosting, lee2025full, lee2025selection and nair2025diversifying compute e-values for (non-)selection events and plug these into the e-BH procedure to control the false discovery rate. gauthier2025values averages conformal e-values over ambiguous ground truths, where observations may carry multiple valid labels simultaneously. Moreover, gauthier2025values also considers e-values to select a prediction set with a fixed number of elements. Our work differs from this line of work, as our intention is to directly report the e-values as output in the form of a fuzzy prediction set, rather than using them as an intermediate step to obtain classical prediction sets.
In the early version, vovk2024conformal was also the first to propose using conformal e-values of the form
$\frac{T(Z_{n+1})}{\frac{1}{n+1}\sum_{i=1}^{n+1}T(Z_{i})},$ (3)
where $T$ is a non-negative statistic. This form was later independently rediscovered in the first arXiv version of wang2022false, koning2023post and balinsky2024enhancing, and proven to be the only admissible type of e-values for conformal prediction by koning2023post. A key issue of the form (3) is that there is little guidance on how to appropriately select this statistic $T$ and gauthier2025values find the outcomes to be highly sensitive to this choice. We provide guidance by introducing a general expected-utility framework to conformal prediction, which enables users to quantify their preference for different degrees of evidence, and allows them to specify a measure $\mu$ that determines the most important regions of the outcome space. This nests classical conformal prediction as a special case for a particular choice of the utility function, as well as the setting studied by vovk2024conformal for a logarithmic choice of utility function and counting measure $\mu$ in classification problems. In our application, we illustrate that log-utility is not always appropriate in the conformal prediction context. This is not surprising, as log-utility is generally considered in long-run sequential data-collection settings instead of the single-batch setting we consider here.
3 Prediction sets
@@
Let $\mathcal{Z}$ denote the sample space of a single observation, and $\mathcal{Z}^{n+1}$ the joint sample space of $(n+1)$ observations. [^1] This is easily generalized to more general joint sample spaces, but we stick to copies of the same sample space for the sake of exposition. A prediction set is classically defined as a set $C_{\alpha}^{Z^{n}}$ with the level $\alpha>0$ coverage guarantee
@@
where $\mathcal{P}$ is some model (collection of distributions) on $\mathcal{Z}^{n+1}$. This means the prediction set $C_{\alpha}^{Z^{n}}$ will cover $Z_{n+1}$ with at least $1-\alpha$ probability. We say that such a prediction set is valid for the model $\mathcal{P}$. Notice that the probability in (4) is over both $Z^{n}$ and $Z_{n+1}$, so that the coverage guarantee is marginal over the entire tuple $Z^{n+1}$.
Conformal prediction is the special case where the model $\mathcal{P}$ is the collection $\mathcal{P}^{\textnormal{exch.}}$ of exchangeable distributions on $\mathcal{Z}^{n+1}$, but we delay specializing to exchangeability in Section 5. Moreover, we postpone the inclusion of covariates to Section 7 as their inclusion considerably clutters the notation and is not relevant for the discussion here.
3.1 Slicing a prediction set for $Z^{n+1}$
Figure 3: Abstract illustration of the sample space $\mathcal{Z}^{n+1}$, where the horizontal dimension represents $\mathcal{Z}^{n}$ and the vertical dimension $\mathcal{Z}_{n+1}$. The prediction set $C_{\alpha}$ for $Z^{n+1}$ is shaded white, and its complement $\mathcal{Z}^{n+1}\setminus C_{\alpha}$ grey. The prediction set $C_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ appears as a $Z^{n}$-dependent random slice of $C_{\alpha}$.
The first key insight is that the construction of a prediction set $C_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ is equivalent to constructing a prediction set $C_{\alpha}$ for $Z^{n+1}$. Indeed, the prediction set $C_{\alpha}^{z^{n}}$ can be interpreted as a ‘slice’ of a prediction set $C_{\alpha}$:
$C_{\alpha}^{z^{n}}=\{z_{n+1}\in\mathcal{Z}:(z^{n},z_{n+1})\in C_{\alpha}\}.$ (5)
We illustrate this in Figure 3, where union of the slices $\{C_{\alpha}^{z^{n}}\}_{z^{n}\in\mathcal{Z}^{n}}$ forms the prediction set $C_{\alpha}$. Here, we stress that the prediction set $C_{\alpha}$ is deterministic: the randomness in $C_{\alpha}^{Z^{n}}$ is purely due to the random realization of the point $Z^{n}$ at which it is ‘sliced’.
The validity of $C_{\alpha}$ as a prediction set for $Z^{n+1}$ under the model $\mathcal{P}$ is equivalent to the validity of $C_{\alpha}^{Z^{n}}$ in the sense of (4). Indeed, by (5) we have $\mathbb{I}\{Z^{n+1}\in C_{\alpha}\}=\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}$, so that
$P(Z^{n+1}\in C_{\alpha})=P(Z_{n+1}\in C_{\alpha}^{Z^{n}}),$
for any $P$. As a consequence, we may focus our attention to constructing prediction sets $C_{\alpha}$ for $Z^{n+1}$.
Remark 1 (Slicing on observables).
These arguments may be extended to slicing based on any part or statistic of $Z^{n+1}$ that is observed. We illustrate this in the covariates setting described in Section 7 where $Z_{n+1}=(X_{n+1},Y_{n+1})$ and we observe $X_{n+1}$ alongside $Z^{n}$, to construct a prediction set for $Y_{n+1}$. There, we still construct a prediction set $C_{\alpha}$ for $Z^{n+1}$ but then slice it at $(Z^{n},X_{n+1})$.
3.2 Prediction as a testing problem
A second insight is that a prediction set $C_{\alpha}$ is equivalent to a hypothesis test. This is a powerful observation, because the construction of hypothesis tests is one of the most deeply studied topics in statistics.
Recall from the introduction that we may view a prediction set $C_{\alpha}$ for $\mathcal{P}$ as a map $C_{\alpha}:\mathcal{Z}^{n+1}\to\{0,1\}$, which returns an inclusion (0) or exclusion (1) for each point $z^{n+1}\in\mathcal{Z}^{n+1}$. We choose to interpret this map $C_{\alpha}$ as a test, where 0 represents a non-rejection and 1 a rejection. Coupling this to Figure 3, the white region corresponds to the non-rejection region of this test and the grey region to its rejection region.
The prediction set $C_{\alpha}$ (and by extension $C_{\alpha}^{Z^{n}}$) is valid for the model $\mathcal{P}$ if and only if it is a valid test for the ‘hypothesis’ $\mathcal{P}$, since
$P(Z_{n+1}\not\in C_{\alpha}^{Z^{n}})=P(Z^{n+1}\not\in C_{\alpha})=\mathrm{E}^{P}[C_{\alpha}],$
for every $P\in\mathcal{P}$. A consequence is that any test that is valid for $\mathcal{P}$ automatically produces a valid prediction set for the model $\mathcal{P}$, and vice-versa.
Given the importance of this observation, we capture it in Theorem 1 for future reference.